# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v1

@@ 5.2 FPGA implementation of the proposed model
The proposed model consists of three types of blocks: ODEBlock, MHSABlock, and DSBlock. In our implementation, all layers in each block are parallelized along the output channel dimension by utilizing multiple DSP blocks. We describe the implementation of ODEBlock and MHSABlock in the following.
@@ 5.2.1 Implementation of ODEBlock
The number of iterations for each ODEBlock is specified through respective control registers. ODEBlock keeps track of a time variable $t$ ($t_{j}=t_{0}+jh$ in Eq. (17)) during iterations ($j=0,\ldots,C-1$). The module “Add time” in Table 1 is to insert an additional channel filled with a time variable $t$ to the layer input; thus, this increases the number of output channels by one. In addition to on-chip buffers for weight parameters, extra buffers are allocated to retain the input to the block for skip connection and the output of each layer.
5.2.2 Implementation of MHSA
@@
MHSA (Fig. 4) is executed in the following sequence: (1) Three weights $\mathbf{W}_{k}$, $\mathbf{W}_{q}$, and $\mathbf{W}_{v}$, along with pre-trained $\mathbf{R}_{w}$ and $\mathbf{R}_{h}$ for positional encoding, are transferred to the on-chip buffers. Then, (2) $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{QK}^{\top}$ are calculated sequentially for input $\mathbf{X}$. (3) For positional encoding, $\mathbf{R}=\mathbf{1}_{w}^{\top}\mathbf{R}_{h}+\mathbf{1}_{h}\mathbf{R}_{w}$ and $\mathbf{QR}^{\top}$ are calculated sequentially. (4) $\mathbf{QK}^{\top}$ is added to the positional encoding $\mathbf{QR}^{\top}$ (i.e., $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$). Afterward, (5) $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$ is divided by a scaling factor $\sqrt{D_{h}}$. (6) ReLU is applied for $\frac{\mathbf{Q}\mathbf{K}^{\top}+\mathbf{Q}\mathbf{R}^{\top}}{\sqrt{D}_{h}}$ to acquire an attention $\mathbf{A}$. Subsequently, (7) $\mathbf{V}$ is calculated for input $\mathbf{X}$. (8) A dot product of $\mathbf{A}$ and $\mathbf{V}$ is calculated. Finally, (9) LayerNorm is applied for the output of MHSA.
@@ 5.3 Quantization of layers
Furthermore, the design performs an LUT-based quantization as described in Sec. 3.3. The accuracy with respect to different quantization levels is evaluated in Sec. 6.2.3. Except the first and last layers in the model, each fully-connected/convolution layer has two learnable LUTs for weights and activations. The LUTs for weights are not necessary during inference, and thus only $n$-bit quantized weights are stored on-chip, which are computed by a host CPU at initialization. On the other hand, the LUTs for activations are transferred to the PL part, since the activations are quantized during inference. In addition, the scaling parameters $s_{a}$ and $s_{w}$ are transferred. The quantization of a single layer introduces additional $22^{n}K+$ 2 parameters for one LUT and two scaling parameters, which increase the on-chip memory footprint by $(n\cdot 2^{n}K+32\cdot 2)$ bits. On the other hand, the memory consumption for weights reduces from $32N_{p}$ to $nN_{p}$ bits, where $N_{p}$ denotes the number of weight parameters in a layer. We assume $K=9$ as in [26]. In this case, the following inequality should be satisfied for the quantization to be effective.
@@
We can derive that $N_{p}>769$ when $n=8$ and $N_{p}>385$ when $n=4$. In our model, the quantization is applied to depth-wise/point-wise convolution layers present in ODEBlock1-2 and MHSABlock. These convolution layers have 585 to 16,448 parameters and thus Eq. (30) holds in most cases except the smallest layer. The memory reduction rate is calculated as follows:
@@
For instance, the point-wise convolution layer in MHSABlock, which is the largest in the model, achieves 3.51$\times$ and 6.99$\times$ reduction when $n=4$ and $8$, respectively. In typical cases where $N_{p}\gg n$, Eq. (31) can be approximated as $32/n$.
@@
Block Output size Input/output channels Architecture
Pre-processing $24\times 24$ 3 (RGB) $\rightarrow$ 64 $7\times 7$ Conv
@@ 6 Evaluations
We compare the proposed model with ODENet [27], BoTNet [7], ViT [2], and lightweight CNN-based architectures such as ResNet18, EfficientNet [33], EfficientNetV2 [34], MobileNet V2 [35], and MobileNet V3 [36] in terms of parameter size and accuracy. Furthermore, we conduct ablation studies to explore the trade-off between accuracy and resource utilization. We investigate the effectiveness of the newly added MHSABlock and the impact of LUT-based quantization on accuracy.
@@ 6.1.1 STL10 dataset
The proposed model is a combination of CNN and attention mechanism. Since the attention mechanism aims to aggregate global information, it is expected to be beneficial in case of larger input images. As discussed in Sec. 2.1, such a hybrid model can achieve a high accuracy even with small datasets compared to pure attention-based models (e.g., ViT). To demonstrate the benefit of the proposed hybrid model, we compare its performance with ViT-Base [2] with a relatively small STL10 [37] dataset containing 10 object classes. STL10 contains 50k labeled images of $96\times 96$ pixels for training and 8k for testing. While CIFAR-10/100 [12] datasets are widely used for image classification models, they contain smaller images (e.g., $32\times 32$) than STL10.
@@ 6.1.2 Model training
All network models considered here are implemented with Python 3.8.10 and PyTorch 1.12.1. The proposed model is built based on the code used in [27]. In our proposed model, DSC is applied to ODEBlocks while it is not applied to DSBlocks. In addition, by applying DSC to DSBlock1, DSBlock2, and both of them, we build three variant models, referred to as “DS1DSC”, “DS2DSC”, and “DS12DSC”, respectively. In this paper, we used the same hyperparameters for training except the batch size; specifically, the batch size is set to 5 for the proposed model, and 128 for the others. The number of epochs is set to 310, and SGD is used as an optimizer. We set the weight decay to $10^{-4}$ and the momentum to 0.9. The cosine annealing with warm restart is used as a learning rate scheduler. The initial learning rate is set to 0.1, and the minimum learning rate is set to $10^{-4}$. For data augmentation, the training images were randomly flipped, erased with a probability of 0.5, and jittered.
Table 2: Parameter size of the proposed and counterpart models
Model Parameter size MHSA
ResNet50 23.52M
BoTNet50 [7] 18.89M
ODENet [27] 1.29M
Proposed model 1.26M
Proposed model (DS1DSC) 1.08M
Proposed model (DS2DSC) 0.51M
Proposed model (DS12DSC) 0.33M
ViT-Base [2] 78.22M
ResNet18 14.19M
EfficientNetb0 [33] 5.33M
EfficientNetv2s [34] 21.61M
MobileNet V2 [35] 3.54M
MobileNet V3 [36] 2.56M
@@ 6.2 Evaluation results
6.2.1 Parameter size
Table 2 shows the parameter size of the proposed and counterpart models with and without MHSA mechanism. As shown, the proposed model uses the smallest number of parameters among these lightweight models, thanks to the benefits of Neural ODE. BoTNet50 has 1.24$\times$ less parameters than ResNet50, because convolutions in the last three ResBlocks are replaced by MHSA. Our proposed model achieves 14.96$\times$ reduction from BoTNet50 by combining MHSA and Neural ODE; as described in Sec. 3, the use of Neural ODE allows to compress $C$ ResBlocks into one ODEBlock, yielding $C$-fold reduction. In our case, $C$, i.e., the number of ODE solver iterations, is set to 10. By employing DSC instead of the original convolution in DSBlocks, the parameter size can be further reduced by up to 3.86$\times$ (see DS1DSC, DS2DSC, and DS12DSC in Table 2). Compared to our proposed model, ViT-Base is 61.96$\times$ larger, making it unsuitable for resource-limited devices due to its high computational cost. While EfficientNet and MobileNet are designed with the deployment on mobile devices in mind, they still have 2.02–17.12$\times$ more parameters than ours, and also do not benefit from the attention mechanism.
@@ 6.2.2 Classification accuracy
Table 3 presents the top-1 accuracy of each model with STL10 dataset. BoTNet shows 2.4% higher accuracy than ResNet50, owing to the attention blocks placed at the end. Analogous to this, our proposed model achieves 1.04% accuracy improvement over ODENet. The proposed model which simplifies the architecture of BoTNet by introducing the concept of Neural ODE leads to a similar improvement on accuracy. While our proposed model has 14.96$\times$ less parameters than BoTNet, it offers comparable accuracy and hence greatly improves the trade-off between accuracy and computational cost. In addition, it achieves 2.89–7.72% better accuracy than EfficientNet and MobileNet series. Fig. 6 visualizes the top-1 accuracy and parameter size of different models. The proposed model is in the top-left most corner, highlighting the better accuracy-computation trade-off. Due to the scaling laws between the number of parameters and accuracy in AI models, models positioned in the upper left of the diagram tend to offer a favorable trade-off.

### v2

@@ 5.2 FPGA implementation of the proposed model
The proposed model consists of three types of blocks: ODEBlock, DSBlock, and MHSABlock. The proposed model is designed with C/C++ with Vitis HLS. The proposed accelerator is implemented as a stream of blocks, as shown in Fig. 6. That is, input samples go through separated blocks, such as ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock. All layers in each block are parallelized along the output channel dimension by utilizing multiple DSP blocks. We describe the implementation of ODEBlock, DSBlock, and MHSABlock in the following.
@@ 5.2.1 Implementation of ODEBlock
Fig. 7 illustrates the ODEBlock implementation. Since the residual connection of ResBlocks is replaced with ODEBlock, the shortcut connection is not included in the ODEBlock. In the ODEBlock, a numerical computation to solve the ordinary differential equation in Eq. (17) is carried out as described in Sec. 3.2. As shown in Table 1, ODEBlocks consist of several modules, such as Add time, DSC, BatchNorm, and ReLU. The weight buffer size for each ODEBlock is determined by DSC and BatchNorm (Add time and ReLU do not have parameters). DSC consists of point-wise and depth-wise convolution layers; thus, the weight buffer size for each DSC is $Ch_{in}\times K\times K+Ch_{out}\times Ch_{in}$. BatchNorm requires two parameters precomputed by software for each channel; thus, the weight buffer size for each BatchNorm is $2\times Ch$. Assuming model sizes in Table 1, the weight buffer sizes for ODEBlock1 and ODEBlock2 are 9,618 and 35,858, respectively.
@@
As described in Sec. 3.2, the number of iterations in ODEBlock is given as a hyperparameter so that we can determine the time interval for the integration. The number of iterations for each ODEBlock is specified through respective control registers. ODEBlock keeps track of a time variable $t$ ($t_{j}=t_{0}+jh$ in Eq. (17)) during iterations ($j=0,\ldots,C-1$). The module “Add time” in Table 1 is to insert an additional channel filled with a time variable $t$ to the layer input; thus, this increases the number of output channels by one. In addition to on-chip buffers for weight parameters, extra buffers are allocated to retain the input to the block for skip connection and the output of each layer.
5.2.2 Implementation of DSBlock
Fig. 8 illustrates the DSBlock implementation. The main purpose of DSBlock is to reduce the feature map size by half in the model. Thus, DSBlocks are not iterated unlike ODEBlocks, and Neural ODE is not applied to the DSBlock in our implementation. DSBlocks consist of several modules, such as convolution, BatchNorm, and ReLU. The weight buffer size for each DSBlock is determined by the convolution and BatchNorm. The weight buffer size for each convolution layer is $Ch_{out}\times Ch_{in}\times K\times K$. The weight buffer size for each BatchNorm is $2\times Ch$ as well as ODEBlock. Assuming model sizes in Table 1, the weight buffer sizes for DSBlock1 and DSBlock2 are 237,047 and 925,943, respectively.
5.2.3 Implementation of MHSABlock
@@
Fig. 9 illustrates the MHSABlock implementation. As well as the ODEBlock, the shortcut connection is not included in the MHSABlock. In the MHSABlock, a spatial convolution layer of the original ResNet building block is replaced with MHSA. MHSA (Fig. 4) is executed in the following sequence: (1) Three weights $\mathbf{W}_{k}$, $\mathbf{W}_{q}$, and $\mathbf{W}_{v}$, along with pre-trained $\mathbf{R}_{w}$ and $\mathbf{R}_{h}$ for positional encoding, are transferred to the on-chip buffers. Then, (2) $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{QK}^{\top}$ are calculated sequentially for input $\mathbf{X}$. (3) For positional encoding, $\mathbf{R}=\mathbf{1}_{w}^{\top}\mathbf{R}_{h}+\mathbf{1}_{h}\mathbf{R}_{w}$ and $\mathbf{QR}^{\top}$ are calculated sequentially. (4) $\mathbf{QK}^{\top}$ is added to the positional encoding $\mathbf{QR}^{\top}$ (i.e., $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$). Afterward, (5) $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$ is divided by a scaling factor $\sqrt{D_{h}}$. (6) ReLU is applied for $\frac{\mathbf{Q}\mathbf{K}^{\top}+\mathbf{Q}\mathbf{R}^{\top}}{\sqrt{D}_{h}}$ to acquire an attention $\mathbf{A}$. Subsequently, (7) $\mathbf{V}$ is calculated for input $\mathbf{X}$. (8) A dot product of $\mathbf{A}$ and $\mathbf{V}$ is calculated. Finally, (9) LayerNorm is applied for the output of MHSA.
The MHSABlock consists of multiple modules, such as point-wise convolution, BatchNorm, and MHSA. The MHSA has weight parameters $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. In addition, those for the positional encoding and LayerNorm are needed. The weight buffer size for $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ is $3\times Ch\times Ch$ in total. That for the positional encoding is $Ch\times H\times W$. LayerNorm also requires two parameters precomputed by software for each feature; thus, the weight buffer size for the LayerNorm is $2\times Ch\times H\times W$. Assuming model sizes in Table 1, the weight buffer size for MHSABlock is 64,433.
Figure 7: ODEBlock implementation
Figure 8: DSBlock implementation
Figure 9: MHSABlock implementation
@@ 5.3 Quantization of layers
Furthermore, the design performs an LUT-based quantization as described in Sec. 3.3. The accuracy with respect to different quantization levels is evaluated in Sec. 6.2.3. Except the first and last layers in the proposed model, each fully-connected/convolution layer has two learnable LUTs for weights and activations. The LUTs for weights are not necessary during inference, and thus only $n$-bit quantized weights are stored on-chip, which are computed by a host CPU at initialization. On the other hand, the LUTs for activations are transferred to the PL part, since the activations are quantized during inference. In addition, the scaling parameters $s_{a}$ and $s_{w}$ are transferred. The quantization of a single layer introduces additional $22^{n}K+$ 2 parameters for one LUT and two scaling parameters, which increase the on-chip memory footprint by $(n\cdot 2^{n}K+32\cdot 2)$ bits. On the other hand, the memory consumption for weights reduces from $32N_{p}$ to $nN_{p}$ bits, where $N_{p}$ denotes the number of weight parameters in a layer. $K$ is set to 9 as in [32]. In this case, the following inequality should be satisfied for the quantization to be effective.
@@
$N_{p}>769$ holds when $n=8$ and $N_{p}>385$ when $n=4$. In the proposed model, the quantization is applied to depth-wise/point-wise convolution layers present in ODEBlock1-2 and MHSABlock. These convolution layers have 585 to 16,448 parameters and thus Eq. (30) holds in most cases except the smallest layer. The memory reduction rate is calculated as follows:
@@
For instance, the point-wise convolution layer in MHSABlock, which is the largest in the proposed model, achieves 3.51$\times$ and 6.99$\times$ reduction when $n=4$ and $8$, respectively. In typical cases where $N_{p}\gg n$, Eq. (31) can be approximated as $32/n$.
@@
In our FPGA implementation, the quantization is selectively applied to some specific blocks by considering the accuracy and parameter size. In Sec. 6.2.3, ablation studies to investigate the quantization will be conducted and show that the model where the quantization is applied to DSBlocks and MHSABlock (denoted as “Block3” in Table 5) strikes a good balance between the accuracy and parameter size. Based on this result, 4-bit or 8-bit quantization is applied to DSBlocks and MHSABlock to implement dedicated IP cores.
5.4 Degree of parallelism in each layer
In our FPGA implementation, all the weight parameters processed with the PL part are stored in on-chip buffers; thus, the main bottleneck in execution time lies in the computation rather than the data transfer. Thus, parallelization is applied to all the layers in Table 1 to speed up the computation as much as the FPGA resources such as BRAM and DSP allow. Based on simulation results with Vitis HLS, a higher degree of parallelism is applied to layers with longer execution time. Table 2 shows the degree of parallelism and the number of cycles for each layer. In this table, SC, DW, PW, and BNReLU stand for shortcut connection, depth-wise, point-wise, and BatchNorm and ReLU layers, respectively. Note that the same degree of parallelism is applied for both the 4-bit and 8-bit quantized implementations.
@@ 5.4 Degree of parallelism in each layer
Block Output In/Out Ch Layer
Pre-processing $24\times 24$ RGB $\rightarrow$ 64 $7\times 7$ Conv
@@ 64 $\rightarrow$ 64 BatchNorm
128 $\rightarrow$ 128 BatchNorm
64 $\rightarrow$ 128 Conv, stride 2 (short cut)
@@
128 $\rightarrow$ 256 Conv, stride 2 (short cut)
256 $\rightarrow$ 256 BatchNorm
@@ 65 $\rightarrow$ 256 $1\times 1$ Conv
Table 2: Degree of parallelism and execution cycles
Block Layer Parallelism Latency (cycles)
ODEBlock1 Add time + DWConv 16 54,721
PW Conv 64 54,720
BNReLU 2 18,439
DSBlock1 $3\times 3$ Conv, stride 2 16 714,251
$3\times 3$ Conv, stride 1 32 688,331
$3\times 3$ Conv, stride 2 (SC) 8 781,483
BNReLU 2 9,223
ODEBlock2 Add time + DW Conv 16 24,625
PW Conv 64 37,160
BNReLU 2 9,223
DSBlock2 $3\times 3$ Conv, stride 2 16 688,907
$3\times 3$ Conv, stride 1 32 676,235
$3\times 3$ Conv, stride 2 (SC) 8 759,627
BNReLU 2 4,615
MHSABlock Add time + $1\times 1$ Conv 16 46,245
BNReLU 2 1,159
MHSA 16 269,448
Add time + $1\times 1$ Conv 16 39,763
@@ 6 Evaluations
The proposed model is evaluated in comparison with ODENet [33], BoTNet [8], ViT [2], and lightweight CNN-based architectures such as ResNet18[11], EfficientNet [39], EfficientNetV2 [40], MobileNet V2 [41], and MobileNet V3 [42] in terms of parameter size and accuracy. Furthermore, we conduct ablation studies to explore the trade-off between accuracy and resource utilization. We investigate the effectiveness of the newly added MHSABlock and the impact of LUT-based quantization on accuracy.
@@ 6.1.1 STL10 dataset
The proposed model is a combination of CNN and attention mechanism. Since the attention mechanism aims to aggregate global information, it is expected to be beneficial in case of larger input images. As discussed in Sec. 2.1, such a hybrid model can achieve a high accuracy even with small datasets compared to pure attention-based models (e.g., ViT). To demonstrate the benefit of the proposed hybrid model, its performance is compared with ViT-Base [2] with a relatively small STL10 [43] dataset containing 10 object classes. STL10 contains 50k labeled images of $96\times 96$ pixels for training and 8k for testing. While CIFAR-10/100 [14] datasets are widely used for image classification models, they contain smaller images (e.g., $32\times 32$) than STL10.
@@ 6.1.2 Model training
All network models considered here are implemented with Python 3.8.10 and PyTorch 1.12.1. The proposed model is built based on the code used in [33]. In the proposed model, DSC is applied to ODEBlocks while it is not applied to DSBlocks. In addition, by applying DSC to DSBlock1, DSBlock2, and both of them, we build three variant models, referred to as “DS1DSC”, “DS2DSC”, and “DS12DSC”, respectively. In this paper, the same hyperparameters are used for training except the batch size; specifically, the batch size is set to 5 for the proposed model, and 128 for the others. The number of epochs is set to 310, and SGD is used as an optimizer. The weight decay is set to $10^{-4}$ and the momentum to 0.9. The cosine annealing with warm restart is used as a learning rate scheduler. The initial learning rate is set to 0.1, and the minimum learning rate is set to $10^{-4}$. For data augmentation, the training images were randomly flipped, erased with a probability of 0.5, and jittered.
Table 3: Parameter size and GFLOPs of the proposed and counterpart models
Model Parameter size GFLOPs MHSA
ResNet50 [11] 25.56M 0.76
BoTNet50 [8] 20.86M 0.74
ODENet [33] 1.29M 0.22
Proposed model 1.26M 0.21
Proposed model (DS1DSC) 1.08M 0.19
Proposed model (DS2DSC) 0.51M 0.19
Proposed model (DS12DSC) 0.33M 0.16
ViT-Base [2] 78.22M 15.46
ResNet18 11.69M 0.34
EfficientNetb0 [39] 5.33M 0.07
EfficientNetv2s [40] 21.61M 0.53
MobileNet V2 [41] 3.54M 0.06
MobileNet V3 [42] 2.56M 0.01
@@ 6.2 Evaluation results
6.2.1 Parameter size and computational cost
Table 3 shows the parameter size and GFLOPs (giga floating-point operations) of the proposed and counterpart models with and without MHSA mechanism. As shown, the proposed model uses the smallest number of parameters among these lightweight models, thanks to the benefits of Neural ODE. BoTNet50 has 1.24$\times$ less parameters than ResNet50, because convolutions in the last three ResBlocks are replaced by MHSA. The proposed model achieves 14.96$\times$ reduction from BoTNet50 by combining MHSA and Neural ODE; as described in Sec. 3, the use of Neural ODE allows to compress $C$ ResBlocks into one ODEBlock, yielding $C$-fold reduction. In this case, $C$, i.e., the number of ODE solver iterations, is set to 10. By this model compression, the proposed model requires a high computational demand relative to parameter size. From the table, it can be observed that GFLOPs per million parameters in ODENet and the proposed model (0.17–0.48) are mostly larger compared to the counterpart models (0.0039–0.20). These models are thus well-suited for hardware accelerators in edge devices with limited memory capacity.
By employing DSC instead of the original convolution in DSBlocks, the parameter size can be further reduced by up to 3.86$\times$ (see DS1DSC, DS2DSC, and DS12DSC in Table 3). Compared to the proposed model, ViT-Base is 61.96$\times$ larger, making it unsuitable for resource-limited devices due to its high computational cost. While EfficientNet and MobileNet are designed with the deployment on mobile devices in mind, they still have 2.02–17.12$\times$ more parameters than the proposed model, and also do not benefit from the attention mechanism.
@@ 6.2.2 Classification accuracy
Table 4 presents the top-1 accuracy of each model with STL10 dataset. BoTNet shows 2.4% higher accuracy than ResNet50, owing to the attention blocks placed at the end. Analogous to this, the proposed model achieves 1.04% accuracy improvement over ODENet. The proposed model which simplifies the architecture of BoTNet by introducing the concept of Neural ODE leads to a similar improvement on accuracy. While the proposed model has 14.96$\times$ less parameters than BoTNet, it offers comparable accuracy and hence greatly improves the trade-off between accuracy and computational cost. In addition, it achieves 2.89–7.72% better accuracy than EfficientNet and MobileNet series. Fig. 10 visualizes the top-1 accuracy and parameter size of different models. The proposed model is in the top-left most corner, highlighting the better accuracy-computation trade-off. Due to the scaling laws between the number of parameters and accuracy in AI models, models positioned in the upper left of the diagram tend to offer a favorable trade-off.
Table 5: Best accuracy of each combination of quantized blocks
Learning rate Model 8-bit acc (%) 4-bit acc (%)
[5e-3,1e-5] Pre-trained 80.15 80.15
(a) [5e-3,1e-5] DS only 81.29 80.06
Block1 76.13 73.21
Block2 75.81 76.51
Block3 81.56 80.78
Block1+2 62.40 65.01
Block1+3 71.76 75.20
Block2+3 77.11 76.21
(b) [5e-4,1e-5] DS only 81.46 80.38
Block1 77.61 78.35
Block2 77.61 76.31
Block3 80.40 79.83
Block1+2 72.30 64.71
Block1+3 76.30 77.15
Block2+3 77.38 78.30
(c) [1e-4,1e-5] DS only 80.97 80.23
Block1 76.59 76.98
Block2 75.94 76.45
Block3 79.10 80.53
Block1+2 66.19 65.6
Block1+3 73.70 75.24
Block2+3 72.26 76.10